Reports of 24772
VAC (Valves & Pumps)
nobuhiro.kimura - 15:55 Friday 29 September 2023 (26985) Print this report
Operation record of the gate valve between X-arm and EXA (September 29)

This is operation record of the gate valve (GVetmx) between X-arm and EXA.
First, residual gas measurements were started to monitor changes in residual gas in the EXC before and after valve operation.

10:40 GVetmx is "Close" position

          Total pressure in EXC: 1.92 x 10^-5 Pa

          H2O: 1.21 x 10^-5 Pa 

          Pressure at X-27: 1.1 x 10^-5 Pa

          Pressure at EXA: 8.5 x 10^-6 Pa

10:42 GVetmx is to "  Open"  position

10:43 Total pressure in EXC: 1.92 x 10^-5 Pa

          H2O: 1.17 x 10^-5 Pa 

 

10:46 Total pressure in EXC: 1.91 x 10^-5 Pa

          H2O: 1.18 x 10^-5 Pa 

          Pressure at EXA: 9.0 x 10^-6 Pa

10:50 Total pressure in EXC: 1.92 x 10^-5 Pa

          H2O: 1.21 x 10^-5 Pa 

          Pressure at EXA: 8.9 x 10^-6 Pa

10:55 Total pressure in EXC: 1.93 x 10^-5 Pa

          H2O: 1.21 x 10^-5 Pa 

          Pressure at EXA: 8.9 x 10^-6 Pa

          Pressure at X-27: 1.1 x 10^-5 Pa         

From these results, it was concluded that the pressure changes before and after opening and closing the GVetmx were minimal.     

        

   

Images attached to this report
DGS (General)
takafumi.ushiba - 15:45 Friday 29 September 2023 (26982) Print this report
New AA installation at OMC0 rack

[Ikeda, Ushiba]

We installed new AA filter (S2214373) on OMC0 rack and change cabling for solving channel confricts.
I first thought all channels on the new AA filter can pass 24kHz signals but I noticed it was wrong (7th and 8th Dsub port had  nominal cutoff filters).
So, I changed the cabling as follows and asked Ikeda-san to adjust the models to the new channel assign.

OMC0 ADC2:
ch17-20: OMMT2T QPDA1 SEG1-4
ch21-24: OMMT2T QPDA2 SEG1-4

OMC0 ADC3:
ch1-4: OMC beacon WFS1 RF17 SEG1/2 I/Q
ch5-8: OMC beacon WFS1 RF17 SEG3/4 I/Q
ch9-12: OMC beacon WFS2 RF17 SEG1/2 I/Q
ch13-16: OMC beacon WFS2 RF17 SEG3/4 I/Q
ch17-20: Q measurement PD signals (I:17, Q:18)
ch21-24: blank
ch25-28: OMC beacon WFS1 DC SEG1-4
ch29-32: OMC beacon WFS2 DC SEG1-4

Note:

Since the number of channels which needs to pass 24kHz signals are increased from the initial design, it is better to change the AA filter cutoff of ch25-32.
After preparing the spare AA filters that psses 24kHz signals in all channels, the current AA filters for ADC3 should be replaced again.

MIF (General)
satoru.ikeda - 15:41 Friday 29 September 2023 (26986) Print this report
Finesse measurement Yarm 20230929

DATE : 2023/09/29 06:30 UTC
ARM : Y
TEMP ITM : 84.5
TEMP ETM : 84.8
NUM : 10
IFO REFL HWP : 168.0
PSL HWP : 20.0
IMC POWER : 5.6
VALUE : 1297.8
ERROR : 14.5

Measured data is stored to /users/Commissioning/data/Finesse/Yarm/20230929-063010

Images attached to this report
DGS (Test bench)
shoichi.oshino - 15:36 Friday 29 September 2023 (26984) Print this report
Testing ADC
I tested different revisions of ADC on the new test bench.
rev.B can be used without any problems, but the Duotone signal with rev.C seems to take one cycle more than rev.B.
rev.B was fine even if the firmware was updated.
I need to try the new RTPC and new IO chassis combination next.
VAC (Tube X)
nobuhiro.kimura - 15:35 Friday 29 September 2023 (26983) Print this report
Comment to VAC pressure X arm 2790 m increased (26940)

Operation record of the gate valve at X-27 (September 28)
First, the turbomolecular pump at site X-27 was confirmed to be in steady state.
At this time, the gate valve between the turbomolecular pump and the T-shape tube was "Open"
and the gate valve between the X-arm duct and the T-shape tube was "Close".
The pressure in the T-shape tube was 1.5 x 10^-6 Pa.
At 13:41, the gate valve between the X-arm duct and the T-shape tube was opened.
Next, the gate valve between the turbo molecular pump and the T-shape tube was then closed.
The pressure gauge reading changed from 1.5 x 10^-6 Pa to 6.1 x 10^-6 Pa.

DGS (General)
koji.nakagaki - 14:47 Friday 29 September 2023 (26980) Print this report
Fiber cabling at the Xend
[Yoshimura,Yamaguchi,Nakagaki]

We routed fiber cable between the DGS rack and the information panel CTX-131 at the X end.
Images attached to this report
DGS (General)
satoru.ikeda - 14:31 Friday 29 September 2023 (26981) Print this report
k1aso,k1ascbeacon,k1sdfmanage model update

Requested by YamaT-san, Ushiba-san

k1aso:
 OMCREFL_QPDA{1,2}_DC_SEG{1,2,3,4}: adc_3_16-23 -> adc_3_24-31

k1ascbeacon:
 adc3_30_31 => MIR
 adc3_24-31 => Terminate 
 MIR 
  In1: AS_PD1_RF_Q -> AS_PD1_RF_I
  In2: AS_PD1_RF_I -> AS_PD1_RF_Q

Other than above
k1sdfmanage:
 Add channel for IMC Picomotor
 

Non-image files attached to this report
IOO (Laser Bench)
takafumi.ushiba - 14:22 Friday 29 September 2023 (26979) Print this report
Comment to PSL pico motor installation (26965)

We disconnected power supply cables of pico driver and network switch installed yesterday.
A power cable of the driver was disconnected at 10:52:23 JST.
A power cable of the network switch was disconnected at 10:53:37 JST.

Figure 1 shows the trend of temperature of PT01-06.
Left and right T cursors show 10:52:23 and 10:53:37 JST, respectively.

After the right T cursor, the temperature seems changed significantly, so the network switch might affect the cryocon outputs.
However, I'm not sure why the temperature outputs became unstable and somehow recovered yesterday.

Images attached to this comment
LAS (bKAGRA laser)
takafumi.ushiba - 14:07 Friday 29 September 2023 (26978) Print this report
Reduce fiber amp. current

I reduced fiber amp. current from 30.3A to 28.7A so that laser output is 20W at 12:10 JST.

IOO (Laser Bench)
shinji.miyoki - 13:08 Friday 29 September 2023 (26977) Print this report
Comment to PSL pico motor installation (26965)

23 min oscillation seems to vanish after reducing the IMC output (= FBL output ?) from 17W to 15.5W level.

Images attached to this comment
DGS (General)
shoichi.oshino - 12:40 Friday 29 September 2023 (26976) Print this report
Comment to Cleaning network optical fiber (26975)
Yamamoto-san noticed me that he could not see the data.
First, I checked if the frame file was generated.
The latest frame files have been created for both DAQ systems.
Next, I checked to see if the directory was mounted on NDS servers.
I can access each frame file directory from NDS servers.
After that, I requested DAQ_KILL from the guardian again.
After the DAQ was restored, I confirmed that I could view the data with ndscope. 
DGS (General)
shoichi.oshino - 12:39 Friday 29 September 2023 (26975) Print this report
Cleaning network optical fiber
As requested by Ikeda-san, I did network optical fiber cleaning.
This is one of the countermeasures for the IPC glitch problem.
 
[Activity log]
1057 k1dc0 PC side in
1058 k1dc0 PC side out
1059 k1dc0 SW side in
1100 k1dc0 SW side out
1101 k1nds0 PC side
1102 k1nds0 SW side
1104 k1nds1 PC side
1105 k1nds1 SW side
1106 k1fw0 PC side in
1107 k1fw0 SW side in
1108 k1fw1 PC side in
1109 k1fw1 SW side in
1111 k1tw0 PC side
1111 k1tw0 SW side
1113 k1tw1 PC side
1113 k1tw1 SW side
1114 k1bcst0 SW side
1116 k1bcst0 PC side
1118 finished
1121 Requested DAQ_KILL
Comments to this report:
shoichi.oshino - 12:40 Friday 29 September 2023 (26976) Print this report
Yamamoto-san noticed me that he could not see the data.
First, I checked if the frame file was generated.
The latest frame files have been created for both DAQ systems.
Next, I checked to see if the directory was mounted on NDS servers.
I can access each frame file directory from NDS servers.
After that, I requested DAQ_KILL from the guardian again.
After the DAQ was restored, I confirmed that I could view the data with ndscope. 
IOO (Laser Bench)
shinji.miyoki - 12:39 Friday 29 September 2023 (26968) Print this report
Comment to PSL pico motor installation (26965)

After this installation or work

  • some PT thermometer shows the strange change of temperature as Fig.1.
  • PMC out shows strange oscillation as Fig.2. (23 minutes) 

Some stuff on the metal mesh and AC adapter might affect the PT thermometer by heat or electrical noise. Or did you put something whose heat can vary on the optical table? Or did you check a stray light was properly dumped around your touched optics? Or some activities of putting the stuff on the mech affected the connection of wiring for PT01-04?

To check the influence, it is better to remove them or check the difference before and after your work.

In addition, some monitors for PC/servers are better to be turned off.

Images attached to this comment
IOO (Laser Bench)
shinji.miyoki - 10:26 Friday 29 September 2023 (26974) Print this report
Comment to PSL pico motor installation (26965)

Before these activities, the PSL mode changed from Observation mode to Working Mode.

After finishing these works, the PSL mode changed from Working mode to Observation mode.

IOO (Laser Bench)
satoru.ikeda - 8:55 Friday 29 September 2023 (26973) Print this report
Comment to PSL pico motor installation (26965)

2023.09.28Activity log
1315 Entering PSL room
1330 Pico motor driver installed, LAN cable wired
1400 Start operation test of motor driver and rebooter after AC connection
1403 Completed equipment installation
1405 Replace PMC2(M4) and begin testing
1535 Returned to original PMC2(M4)
1540 Exit from PSL room
Times are approximate.

To investigate the cause, the copi motor driver, rebooter, and network switch are all powered on but not connected to the motor, so they can be turned off by entering the PSL room.
 

Non-image files attached to this comment
FCL (Air)
shinji.miyoki - 8:36 Friday 29 September 2023 (26971) Print this report
Comment to IX & EY room temperature (26943)

From 23.5 to 23.0 on 28th Sept.

IOO (Laser Bench)
takaaki.yokozawa - 8:23 Friday 29 September 2023 (26970) Print this report
Comment to PSL pico motor installation (26965)

I may not understand the situation correctly, there are possibility that the pin assignment affect something.
5 pin used in cryocon, as shown in JGWDoc14102.

 

IOO (Laser Bench)
shinji.miyoki - 6:08 Friday 29 September 2023 (26969) Print this report
Comment to PSL pico motor installation (26965)

At present, the PT01~04 output became stable. So this fluctuation comes from the unstability of the cryocon or wiring connection trouble due to yesterday's activity on the metal mesh? Ikeda-san and Yokozawa-kun, do you have any idea?

Actually, the fluctuation is sometimes so fast. It cannot be induced by thermal influence from the mesh area to the optical table.

On the other hand, PMC output oscillation continues.

DetChar (General)
hirotaka.yuzurihara - 23:36 Thursday 28 September 2023 (26956) Print this report
Pastavi update : Added option for SVG format and numpy binary including all the noise budget data

I added several features on Pastavi, based on the comments from commissioners.

  • I added an option to generate a plot in the SVG forma [image]. You can download the generated plot in SVG format from the link [image].
    • The function to make SVG plot is copied from the aligoNB repo and modified to fit the Pastavi code.
    • It takes ~10 seconds to make the SVG plot. So, to relieve stress, I put the option check box.
  • I added the numpy binary file including all the noise budget data [image]. I summarized how to read the noise budget data from the numpy file, in [JGWdoc].
    • This data is the noise budget for O4a generated by Komori-san.
  • I added a link to all the uploaded xml files [image].
Images attached to this report
CAL (General)
hirotaka.yuzurihara - 22:53 Thursday 28 September 2023 (26967) Print this report
TCam photo session 2023/09/28 morning

Yokozawa-san had the TCam photo session before closing the GV of the IXA and EXA this morning (07:57 ~ 8:05 JST). I will cancel the weekly TCam photo session planned on Friday early morning.

  • There is no significant change in the mirror centers from the images taken on 9/26.
  • I found that the blow part of the ITMX image differs from the previous image. [9/26] [9/28] I will check the interferometer condition at that time, tomorrow. 
Images attached to this report
MIF (General)
takafumi.ushiba - 18:38 Thursday 28 September 2023 (26966) Print this report
Comment to Finesse measurement Yarm 230928 (26939)

I measured the Y arm finesse in the different alignments: one is the alignment recorded during PRFPMI lock yesterday and the other is the one that the beam on both ITM and ETM is roughtly centered by using camera.
Figure 1 shows the decay and fiting curve when the mirror alignment was former.
Figure 2 shows the decay and fiting curve when the mirror alignment was latter.

Though there are strange values in both measurement, the finesse value with the latter alignment seems slightly better.
During the measurement, I checked the beam spot on ITMY and it seems shifted several cm horizontally, so this finesse different might come from the clipping of the beam by recoil mass but I'm not so sure.

Anyway, the finesse value measured with the latter alignment is over 1300, finesse degradation seems not happened.

Note:

Somehow, fitting sometimes failed and finese value becomes negative (and absolute value is very large).
So, it might be better to set some limit so that finesse becomes positive.

Images attached to this comment
IOO (Laser Bench)
masayuki.nakano - 17:43 Thursday 28 September 2023 (26965) Print this report
PSL pico motor installation

[Ikeda, Nakano]

We installed the pico driver, rebooter, and network switch for the PMC/IMC remote alignment. Although all of the component has been installed in the PSL, the pico is not installed into the laser path. This was because we could not find a proper mount fitting the pico. We are now trying to find or buy a good mount, then we're gonna install them again.

- The components are located on the head shelf at -X side of the table. 
- We tested the pico itself, and it was working well.
- The channel name of the pico will be PMC and IMC (now it's named POL, but we change it in the tomorrow maintenance). 

Images attached to this report
Comments to this report:
shinji.miyoki - 6:08 Friday 29 September 2023 (26969) Print this report

At present, the PT01~04 output became stable. So this fluctuation comes from the unstability of the cryocon or wiring connection trouble due to yesterday's activity on the metal mesh? Ikeda-san and Yokozawa-kun, do you have any idea?

Actually, the fluctuation is sometimes so fast. It cannot be induced by thermal influence from the mesh area to the optical table.

On the other hand, PMC output oscillation continues.

takaaki.yokozawa - 8:23 Friday 29 September 2023 (26970) Print this report

I may not understand the situation correctly, there are possibility that the pin assignment affect something.
5 pin used in cryocon, as shown in JGWDoc14102.

 

satoru.ikeda - 8:55 Friday 29 September 2023 (26973) Print this report

2023.09.28Activity log
1315 Entering PSL room
1330 Pico motor driver installed, LAN cable wired
1400 Start operation test of motor driver and rebooter after AC connection
1403 Completed equipment installation
1405 Replace PMC2(M4) and begin testing
1535 Returned to original PMC2(M4)
1540 Exit from PSL room
Times are approximate.

To investigate the cause, the copi motor driver, rebooter, and network switch are all powered on but not connected to the motor, so they can be turned off by entering the PSL room.
 

Non-image files attached to this comment
shinji.miyoki - 10:26 Friday 29 September 2023 (26974) Print this report

Before these activities, the PSL mode changed from Observation mode to Working Mode.

After finishing these works, the PSL mode changed from Working mode to Observation mode.

shinji.miyoki - 12:39 Friday 29 September 2023 (26968) Print this report

After this installation or work

  • some PT thermometer shows the strange change of temperature as Fig.1.
  • PMC out shows strange oscillation as Fig.2. (23 minutes) 

Some stuff on the metal mesh and AC adapter might affect the PT thermometer by heat or electrical noise. Or did you put something whose heat can vary on the optical table? Or did you check a stray light was properly dumped around your touched optics? Or some activities of putting the stuff on the mech affected the connection of wiring for PT01-04?

To check the influence, it is better to remove them or check the difference before and after your work.

In addition, some monitors for PC/servers are better to be turned off.

Images attached to this comment
shinji.miyoki - 13:08 Friday 29 September 2023 (26977) Print this report

23 min oscillation seems to vanish after reducing the IMC output (= FBL output ?) from 17W to 15.5W level.

Images attached to this comment
takafumi.ushiba - 14:22 Friday 29 September 2023 (26979) Print this report

We disconnected power supply cables of pico driver and network switch installed yesterday.
A power cable of the driver was disconnected at 10:52:23 JST.
A power cable of the network switch was disconnected at 10:53:37 JST.

Figure 1 shows the trend of temperature of PT01-06.
Left and right T cursors show 10:52:23 and 10:53:37 JST, respectively.

After the right T cursor, the temperature seems changed significantly, so the network switch might affect the cryocon outputs.
However, I'm not sure why the temperature outputs became unstable and somehow recovered yesterday.

Images attached to this comment
PEM (Center)
tatsuki.washimi - 17:09 Thursday 28 September 2023 (26960) Print this report
Test of the new 3-axial seismometer

[YokozaWashimi]

We purchased a new accelerometer, TEAC710Z (3-axial).
We started to test it near the MCF seismometer.

Test for the portable seismometer is also performed.

Images attached to this report
MIF (ASC)
Elenna Capote - 17:06 Thursday 28 September 2023 (26963) Print this report
Dark Noise Analysis of RF QPDs

I measured the dark noise on all the RF QPDs at the REFL, POP and AS ports and quantified how much is ADC noise vs electronics dark noise. I changed the whitening gain of the QPDs to determine if the RF QPDs are limited by ADC noise or electronics dark noise. This work is in preparation for moving to the wavefront sensors for the alignment control, so we can understand the noise limits of the wavefront sensors.

Conclusion: all RF QPDs are using a flat frequency independent whitening/anti-whitening scheme. This is sufficient for the time being at the REFL and POP ports. The AS port QPDs are close to being ADC noise limited, but the current whitening gain is high. We should consider moving to the frequency dependent whitening filters, and probably consider 3 stages of whitening/anti-whitening at the AS port. 2 stages will likely be sufficient for REFL and POP. A frequency dependent whitening/antiwhitening scheme will probably be better when we use these QPDs in loop.

Procedure:

I took the IMC offline and shuttered the IR and green shutters for this test. First, I took spectra of all the QPDs at their nominal whitening gain settings. Then, I set the whitening gain to 1 dB and retook the data.

The first set of data will show the electronics dark noise of the PDs and the ADC noise with the whitening/antiwhitening gain applied. The electronics noise experiences both the whitening gain and antiwhitening gain (so ithe whitening effect cancels out). The ADC noise only experiences the antiwhitening gain. This is the "nominal" dark noise level for these PDs. N1 = n_dark + n_adc*(A.W.)

The second set of data will show the electronics dark noise and ADC noise without any whitening gain. N2 = n_dark + n_adc

To calculate the ADC noise level, I can subtract these two measurements and divide out the effect of the anti-whitening gain.

n_{adc} = \sqrt{\frac{N_2^2 - N_1^2}{1 - AW^2}}

The ADC noise level can then be subtracted from measurement 2 to determine the electronics dark noise level.

When I finished my measurements I made sure to revert all gains properly and reopen the shutters.

I calibrated these measurements from counts to volts with the conversion factor 6.1e-4 V/ct, from a +/-20V, 16 bit ADC.

Results:

I made plots of all these results showing the nominal dark noise level, the ADC noise level, and the current ADC noise level with the whitening gain in place. This is a lot of plots, so I am only attaching a select few that I feel are representative.

POP: The first plot is the result of POP QPD1 RF17 Pitch. All POP results were about the same. Currently there is 21 dB of whitening gain applied for all sensors.

REFL: REFL RF17 is using 21 dB gain for both QPD1 and 2, but RF45 has 24 dB gain for QPD1 and 27 dB gain for QPD2. All QPDs are about the same, except QPD2 RF45 which seems to have slightly lower electronics dark noise than the others. I am attaching The QPD2 RF45 plot, and one other representative REFL plot for comparison

AS: All QPDs are using 36 dB gain at the AS port. Again, all QPDs are about the same, except QPD2 RF28. It looks like the ADC noise is slightly elevated compared to the others. I again am attaching the strange plot and one other representative for comparison.

My next goal is to determine the shot noise level.

Images attached to this report
DGS (General)
shoichi.oshino - 16:43 Thursday 28 September 2023 (26964) Print this report
Comment to Moving k1nds2 to the SK building (26894)
I started to restore the Debian10 test bench system in the SK computer room.
First, I connected IRIG-B and HIB cable to V4 computer, and also connected the optical fiber of timing signal to IO chassis.
I started k1ioo after starting k1boot.
The RT model did not start up, so I restarted the model and it worked.
Search Help
×

Warning

×